How do I find the slope throught (6,1) and perpendicular to y= 3/2x + 1/4
Answer by nerdybill(7384)

How do I find the slope throught (6,1) and perpendicular to y= 3/2x + 1/4
.
Since the slope of:
y= 3/2x + 1/4
is 3/2
.
a line perpendicular to it must be the "negative reciprocal".
That is, our new slope multiplied by 3/2 must be -1:
(3/2)m = -1
3m = -2
m = -2/3
.
Now, if you want the equation of the line passing through (6,1) with a slope of -2/3 plug it into the "point-slope" form:
y - y1 = m(x - x1)
y - 1 = (-2/3)(x - 6)
y - 1 = (-2/3)x + 4
y = (-2/3)x + 5 (this is what they're looking for in "slope-intercept" form)